Circular No. 8717

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 73P/SCHWASSMANN-WACHMANN M. L. Sitko, University of Cincinnati and Space Science Institute; R.

W. Russell, D. K. Lynch, R. Ford, and K. Crawford, The Aerospace Corporation; E. F. Polomski, University of Minnesota; and W. Golisch, D. Griep, and P. Sears, Infrared Telescope Facility (IRTF), NASA, report on observations made of comet 73P’s components ‘B’ and ‘C’ using the IRTF (+ BASS) on May 18 and 19 UT. Both components exhibited a smooth continuum between 3 and 13 μm, on top of which a silicate emission band from 8.5 to 12.2 μm was observed. For component ‘C’, an underlying blackbody, normalized to the continuum fluxes at 7.8 and 12.5 μm, yielded grain temperatures of 295 K on May 18.73, 295 K on May 19.65, and 300 K on May 19.75, with an estimated uncertainty of ± 5 K. For ‘B’, the temperatures were 285 K on May 18.66 and 300 K on May 19.70. These derived temperatures were between 2 and 7 percent higher than that of an equivalent blackbody at the heliocentric distance of the comet. The feature-to-continuum ratios in the silicate band for ‘C’ were 1.25 on May 18.73, 1.25 on May 19.65, and 1.23 on May 19.75. For ‘B’, the ratio was 1.18 for both nights. These values are consistent with those previously reported for ‘B’ observed on Apr. 29.5 (IAUC 8708), but the silicate band in ‘C’ was slightly weaker than on Mar. 17.05 (IAUC 8701). On both nights in May, the spectrum of ‘B’ showed the possible presence of the 3.4-μm organic emission band. The feature was undetected in ‘C’. The measured flux between 10.5 and 11.0 μm, using the 3′′.4 circular entrance aperture of BASS, was 28.7 ± 0.3 Jy (equivalent magnitude m_{N} = 0.37 ± 0.01) on May 18.73. On May 19.76, the flux had dropped to 9.0 ± 0.2 Jy (m_{N} = 1.62 ± 0.02), but brightened to 15.0 ± 0.2 Jy (m_{N} = 1.07 ± 0.01) only 2 hr later. For ‘B’, the flux was 7.7 ± 0.1 Jy (m_{N} = 1.79 ± 0.01) on May 18.66 and 5.5 ± 0.02 Jy (m_{N} = 2.16 ± 0.01) on May 19.70.

COMETS C/2006 G3 AND C/2006 H2 (SOHO) Additional Kreutz sungrazing comets have been found on SOHO web-

site images (cf. IAUC 8714). C/2006 G3 was elongated and diffuse, reaching mag 7.5 at 7.1R_{⊙} on Apr. 14.563 UT, when it had an 84′′ tail. C/2006 H2 was tiny but slightly elongated and too faint for photometry.

Comet 2006 UT α_{2000} δ_{2000} Inst. F MPEC
C/2006 G3 Apr. 14.526 1^{h}38.^{m}8 + 8^{o}38′ C2 HS 2006-K29
C/2006 H2 Apr. 17.493 1 49.0 + 9 40 C2 HS 2006-K29
2006 June 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green